# Characteristic Polynomials and Hypergraph Generating Functions via Heaps of Pieces

### v1

@@ 2.5 The Characteristic Polynomial of a Graph as a sum over Trivial Heaps
Let ~$\widetilde{\mathcal{B}}^{e}(\mathcal{G})$, or shortly ~$\widetilde{\mathcal{B}}^{e}$ be the union of cycles and edgegons of $\mathcal{G}$, containing the edge $e$. (There is a single edgegon.) By Theorem 2.8, we get:
$\dfrac{\widetilde{\phi}_{\mathcal{G}-e}(t)}{\widetilde{\phi}_{\mathcal{G}}(t)}=\dfrac{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}}\setminus\widetilde{\mathcal{B}}^{e})}(-1)^{c(T)}w(T)}{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}})}(-1)^{c(T)}w(T)}=\sum_{\begin{subarray}{c}H\in\mathbbm{h}(\widetilde{\mathcal{B}})\\ \mathcal{M}(H)\subseteq\widetilde{\mathcal{B}}^{e}\end{subarray}}w(H)=\sum_{\begin{subarray}{c}P\in\mathbbm{p}(\widetilde{\mathcal{B}})\\ \mathcal{M}(P)\subseteq\widetilde{\mathcal{B}}^{e}\end{subarray}}w(P)$
@@
$=\sum_{\mathbf{w}\in\mathcal{U}^{e}(\mathcal{G})}t^{-|\mathbf{w}|}=\sum_{d\geq 0}w^{e}_{d}t^{-d}$ by Corollary 2.7 Part 1
By an abuse of notation, let us write $\mathcal{M}(P)$ for the unique piece of a pyramid $P\in\mathbbm{p}(\mathcal{G})$. By combining Theorem 2.11 and Corollary 2.10, we obtain an interesting identity:
@@
$\sum_{P\in\mathbbm{p}(\mathcal{G})}\frac{1}{|P|}t^{-|E(P)|}=\sum_{P\in\mymathbb{p}(\mathcal{G})}\frac{|V(\mathcal{M}(P))|}{|E(P)|}t^{-|E(P)|}$
@@
$\sum_{P\in\mathbbm{p}_{d}(\mathcal{G})}\frac{1}{|P|}=\frac{1}{d}\sum_{P\in\mymathbb{p}_{d}(\mathcal{G})}|V(\mathcal{M}(P))|$
where $\mymathbb{p}_{d}(\mathcal{G})$ is the set of pyramids with $d\geq 0$ edges.
@@ 2.6 An Example: Heaps of Pieces of a Simple Graph
    Trivial Heaps. Using Sagemath ([24]), we calculate
[…]
@@
    $-\sum_{\mathbf{w}\in\mathcal{U}(\mathcal{G})}\frac{1}{|\mathbf{w}|}t^{-|\mathbf{w}|}=\ldots-t^{-4}\left(\frac{1}{4}\cdot 4\cdot 2\cdot 3+\frac{1}{4}\cdot 2\cdot 1\cdot 6+\frac{1}{4}\cdot(1+2+1)\cdot 12\right)-\ldots$
    $-\sum_{P\in\mathbbm{p}(\mathcal{G})}\frac{1}{|P|}w(P)=\ldots-t^{-4}\left(\dfrac{1}{1}\cdot 2\cdot 3+\dfrac{1}{2}\cdot 6+\dfrac{1}{2}\cdot 2\cdot 12\right)-\ldots$
@@ 3.1.2 Rooted stars and Rootings
Let $X=([n],E,\varphi)$ be a Veblen hypergraph. Given an edge $e\in E(X)$, for each $u\in e$, we define the $u$-rooted directed star of $e$, denoted $S_{u}(e)$, as the digraph $T$ with $V(T)=\varphi(e)$ and $E(T)=\{(u,v):v\in\varphi(e)\text{ and }v\neq u\}$.
Let $X=([n],\{e_{1},\ldots,e_{d}\},\varphi)$ be a Veblen hypergraph. Given $X\in\mathcal{V}^{\infty}_{d}$, then a rooting is a $d$-tuple of stars $(S_{v_{1}}(e_{1}),\ldots,S_{v_{d}}(e_{d}))$ such that
@@
From each rooting $R=(S_{v_{1}}(e_{1}),\ldots,S_{v_{d}}(e_{d}))$ of Veblen hypergraph $X$, we obtain a digraph by taking the union of the edges of stars, adding multiplicities when two edges are parallel. Formally, let $D_{R}$ be the digraph with $V(D_{R})=V(X)$, $E(D_{R})=\{(v,w,j):v=v_{j},w\in e_{j}\setminus\{v\}\}$, and $\psi(v,w,j)=(v,w)$ for each $v,w\in[n]$. Note that $m_{D_{R}}(u,v)=|\{j\in[d]:u=v_{j},v\in e_{j}\setminus\{u\}\}|$, for each $u,v\in[n]$, i.e., the number of stars $S_{u}(e_{j})$ in $R$ so that $v\in e_{j}\setminus\{u\}$. We say that $R$ is an Eulerian rooting if $D_{R}$ is Eulerian. Let $\mathfrak{R}(X)$ be the set of Eulerian rootings of $X$.
For a rooting $R\in\mathfrak{R}(X)$, let $i$ be chosen as the root of $r_{i}$-many stars,
@@
For example, if $X$ is the infragraph with $V(X)=[2]$ and $E(X)=\{e_{1},e_{2},e_{3},e_{4}\}$ with $\varphi(e_{j})=\{1,2\}$ for each $j$, then $\mathfrak{R}(X)=\{R\}$, where
@@
since $S_{i}(e_{1})=S_{i}(e_{2})=S_{i}(e_{3})=S_{i}(e_{4})$, for $i=1,2$.
Given a rooting $R=(S_{v_{1}}(e_{1}),\ldots,S_{v_{d}}(e_{d}))$, then rootings obtained by interchanging two stars with the same root and possibly different edges will be considered equivalent to $R$. Formally, we define an equivalence relation $\equiv$ on $\mathfrak{R}(X)$, as the transitive closure of the relations,
$(S_{v_{1}}(e_{1}),\ldots,S_{v_{i}}(e_{i}),S_{v_{i+1}}(e_{i+1})\ldots,S_{v_{d}}(e_{d}))$ $\equiv^{\prime}(S_{v_{1}}(e_{1}),\ldots,S_{v_{i+1}}(e_{i+1}),S_{v_{i}}(e_{i}),\ldots,S_{v_{d}}(e_{d}))$
where $v_{i}=v_{i+1}$ and $11\leq i\leq d-$ 1.
Let $\mathfrak{R}(X)/\!\equiv$ be the set of rootings of $X$, up to the equivalence relation $\equiv$. Instead of $[R]_{\equiv}\in\mathfrak{R}(X)/\!\equiv$, we will shortly write $\mathbf{R}\in\overline{\mathfrak{R}}(X)$. Note that $R\equiv R^{\prime}$ implies $D_{R}=D_{R^{\prime}}$, which allows us to define $D_{\mathbf{R}}:=D_{R}$, where $R\in\mathbf{R}$ is any choice of representative under $\equiv$.
@@ 3.2 Kocay’s Lemma and Alternative Expression of the Characteristic Polynomial
    where $S/\mathord{\sim}=\{[B_{1}]_{\mathord{\sim}},\ldots,[B_{j}]_{\mathord{\sim}}\}$ and $\{B_{1},\ldots,B_{j}\}$ is any choice of representatives of $S/\mathord{\sim}$.
@@ 5 Future Directions
    The rational function $\frac{\phi_{\mathcal{G}-u}(t)}{\phi_{\mathcal{G}}(t)}$ can be used to show the Eigenvalue Interlacing Theorem ([12, Theorem 5.3, p. 71]). Is there a way to generalize this method to hypergraphs?
@@
    In [12, p. 71], the following identity for the characteristic polynomial of a graph $\mathcal{G}$ is given:
    $\phi_{\mathcal{G}}^{\prime}(t)^{2}-\phi_{\mathcal{G}}^{\prime\prime}(t)\phi_{\mathcal{G}}(t)=\sum_{i,j}\phi^{i,j}_{\mathcal{G}}(t)^{2}$
    where $\phi^{i,j}_{\mathcal{G}}(t)$ is the $ij$th entry of the adjugate matrix of $t\cdot I-\mathbb{A}_{\mathcal{G}}$. Is there a version of this identity for hypergraphs?

### v3

@@ 2.5 The Characteristic Polynomial of a Graph as a sum over Trivial Heaps
Let ~$\widetilde{\mathcal{B}}^{e}(G)$, or shortly ~$\widetilde{\mathcal{B}}^{e}$ be the union of cycles and edgegons of $G$, containing the edge $e$. (There is a single edgegon.) By Theorem 2.8, we get:
$\dfrac{\widetilde{\phi}_{G-e}(t)}{\widetilde{\phi}_{G}(t)}=\dfrac{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}}\setminus\widetilde{\mathcal{B}}^{e})}(-1)^{c(T)}w(T)}{\sum_{T\in\mathbbm{t}(\widetilde{\mathcal{B}})}(-1)^{c(T)}w(T)}=\sum_{\begin{subarray}{c}H\in\mathbbm{h}(\widetilde{\mathcal{B}})\\ \mathcal{M}(H)\subseteq\widetilde{\mathcal{B}}^{e}\end{subarray}}w(H)=\sum_{\begin{subarray}{c}P\in\mathbbm{p}(\widetilde{\mathcal{B}})\\ \mathcal{M}(P)\subseteq\widetilde{\mathcal{B}}^{e}\end{subarray}}w(P)$
@@
$=\sum_{\mathbf{w}\in\mathcal{U}^{e}(G)}t^{-|\mathbf{w}|}=\sum_{d\geq 0}w^{e}_{d}t^{-d}$ by Corollary 2.7 Part 1
By an abuse of notation, let us write $\mathcal{M}(P)$ for the unique piece of a pyramid $P\in\mathbbm{p}(G)$. By combining Theorem 2.11 and Corollary 2.10, we obtain an interesting identity:
@@
$\sum_{P\in\mathbbm{p}(G)}\frac{1}{|P|}t^{-|E(P)|}=\sum_{P\in\mymathbb{p}(G)}\frac{|V(\mathcal{M}(P))|}{|E(P)|}t^{-|E(P)|}$
@@
$\sum_{P\in\mathbbm{p}_{d}(G)}\frac{1}{|P|}=\frac{1}{d}\sum_{P\in\mymathbb{p}_{d}(G)}|V(\mathcal{M}(P))|$
where $\mymathbb{p}_{d}(G)$ is the set of pyramids with $d\geq 0$ edges.
@@ 2.6 An Example: Heaps of Pieces of a Simple Graph
    Trivial Heaps. Using Sagemath ([24]), we calculate:
    $\widetilde{\phi}_{G}(t)=1-6t^{-1}-8t^{-2}-3t^{-4}$
@@
    $-\sum_{\mathbf{w}\in\mathcal{U}(G)}\frac{1}{|\mathbf{w}|}t^{-|\mathbf{w}|}=\ldots-t^{-4}\left(\frac{1}{4}\cdot 4\cdot 2\cdot 3+\frac{1}{4}\cdot 2\cdot 1\cdot 6+\frac{1}{4}\cdot(1+2+1)\cdot 12\right)-\ldots$
    $-\sum_{P\in\mathbbm{p}(G)}\frac{1}{|P|}w(P)=\ldots-t^{-4}\left(\dfrac{1}{1}\cdot 2\cdot 3+\dfrac{1}{2}\cdot 6+\dfrac{1}{2}\cdot 2\cdot 12\right)-\ldots$
@@ 3.1.2 Rooted stars and Rootings
Let $X=([n],E,\varphi)\in\mathcal{V}^{\infty}_{d}$ be a Veblen hypergraph with $d\geq 0$ edges. Given an edge $e\in E(X)$, for each $u\in e$, we define the $u$-rooted directed star of $e$, denoted $S_{u}(e)$, as the digraph $T$ with $V(T)=\varphi(e)$ and $E(T)=\{(u,v):v\in\varphi(e)\text{ and }v\neq u\}$. A rooting of $X$ is a $d$-tuple of stars $(S_{v_{1}}(e_{1}),\ldots,S_{v_{d}}(e_{d}))$ such that
@@
From each rooting $R=(S_{v_{1}}(e_{1}),\ldots,S_{v_{d}}(e_{d}))$ of Veblen hypergraph $X$, we obtain a digraph by taking the union of the edges of stars, adding multiplicities when two edges are parallel. Formally, let $D_{R}$ be the digraph with $V(D_{R})=V(X)$, $E(D_{R})=\{(v,w,j):v=v_{j},w\in e_{j}\setminus\{v\}\}$, and $\psi(v,w,j)=(v,w)$ for each $v,w\in[n]$. Note that $m_{D_{R}}(u,v)=|\{j\in[d]:u=v_{j},v\in e_{j}\setminus\{u\}\}|$, for each $u,v\in[n]$, i.e., the number of stars $S_{u}(e_{j})$ in $R$ so that $v\in e_{j}\setminus\{u\}$. We say that $R$ is an Eulerian rooting if $D_{R}$ is Eulerian. Let $\mathfrak{R}(X)$ be the set of Eulerian rootings of $X$. For a rooting $R\in\mathfrak{R}(X)$, let $i$ be chosen as the root of $r_{i}$-many stars,
@@
By the definition of a rooting, it follows that for each pair of parallel edges $e,e^{\prime}$ and each vertex $i\in[n]$, we have $S_{i}(e)=S_{i}(e^{\prime})$. For example, if $X$ is the infragraph with $V(X)=[2]$ and $E(X)=\{e_{1},e_{2},e_{3},e_{4}\}$ with $\varphi(e_{j})=\{1,2\}$ for each $j$, then $\mathfrak{R}(X)=\{R\}$, where
@@
since $S_{i}(e_{1})=S_{i}(e_{2})=S_{i}(e_{3})=S_{i}(e_{4})$, for $i=1,2$. Two rootings are considered equivalent, if one is obtained from the other by interchanging two stars with the same root (and possibly non-parallel edges). Formally, we define an equivalence relation $\equiv$ on $\mathfrak{R}(X)$, as the transitive closure of the relations,
$(S_{v_{1}}(e_{1}),\ldots,S_{v_{i}}(e_{i}),S_{v_{i+1}}(e_{i+1})\ldots,S_{v_{d}}(e_{d}))$ $\equiv(S_{v_{1}}(e_{1}),\ldots,S_{v_{i+1}}(e_{i+1}),S_{v_{i}}(e_{i}),\ldots,S_{v_{d}}(e_{d}))$
where $v_{i}=v_{i+1}$ and $11\leq i\leq d-$ 1. Let $\mathfrak{R}(X)/\!\equiv$ be the set of rootings of $X$, up to the equivalence relation $\equiv$. Instead of $[R]_{\equiv}\in\mathfrak{R}(X)/\!\equiv$, we succinctly write $\mathbf{R}\in\overline{\mathfrak{R}}(X)$. Note that $R\equiv R^{\prime}$ implies $D_{R}\approx D_{R^{\prime}}$. Hence, we the following are well-defined, independent of the choice of representative:
$D_{\mathbf{R}}:=[D_{R}]_{\approx}\text{ and }\mathfrak{C}(D_{\mathbf{R}}):=\mathfrak{C}(D_{R})\text{ for any $R\in\mathbf{R}$.}$
@@ 3.2 Kocay’s Lemma and Alternative Expression of the Characteristic Polynomial
    where $S/\mathord{\sim}=\{[B_{1}]_{\mathord{\sim}},\ldots,[B_{r}]_{\mathord{\sim}}\}$ and $\{B_{1},\ldots,B_{r}\}$ is any choice of representatives of $S/\mathord{\sim}$ and $1\leq r\leq m$.
@@ 5 Future Directions
    The rational function $\frac{\phi_{G-u}(t)}{\phi_{G}(t)}$ can be used to show the Eigenvalue Interlacing Theorem ([12, Theorem 5.3, p. 71]). Is there a way to generalize this method to hypergraphs?
@@
    In [12, p. 71], the following identity for the characteristic polynomial of a graph $G$ is given:
    $\phi_{G}^{\prime}(t)^{2}-\phi_{G}^{\prime\prime}(t)\phi_{G}(t)=\sum_{i,j}\phi^{i,j}_{G}(t)^{2}$
    where $\phi^{i,j}_{G}(t)$ is the $ij$th entry of the adjugate matrix of $t\cdot I-\mathbb{A}_{G}$. Is there a version of this identity for hypergraphs?